[Brahms-l] Shift report 20040320 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Mar 20 2004 - 16:05:15 EST
------------------------------------------------------------
Date: Sat Mar 20 16:05:14 2004, Shift: 08:00-16:00
Supervisor: PS, with: SN
Log book#15 pages 92 - 92

Shift Summary:
------------------------------------------------------------
1. Long store ended at 10.30. Had some very nice beam from 12.

2. 11174   Minimum bias
   11175   Minimum bias
   11176   Non-TPC Pedestal

           tr2     tr3     tr6
   11177   62k     55k     82k
   11178   115k    107k    157k
   11179   115k    122k    175k
   11184   82k     96k     138k
   11185   170k    161k    245k
   11186   100k    108k    162k
   TOTAL   646k    649k    963k 

   11187   In progress. (Will get us above 1M on trigger 6.)

    PS worked on mising 4 planes in the first module
    of T4. He found out that the problem is caused by
    mitake in cabling, namely the D1 and D2 cables  
    (providing connection   between max. board and tdc) are  
    repleaced.Should be fixed during the next access.     

3. T2 tripped, ending run 11184. Well behaved rest of the
   time.
   The measured values of the gas line flow rates were (on
   average) 0.4 above the requested values.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 20 16:05:37 2004

This archive was generated by hypermail 2.1.8 : Sat Mar 20 2004 - 16:06:01 EST